40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Mar 22 16:29:04 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
    Reply  Thu Mar 24 19:07:38 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
       Reply  Sat Mar 26 13:36:09 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT IMG_2468.JPGIMG_2466.JPG
          Reply  Tue Mar 29 23:23:37 2016, rana, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
             Reply  Wed Mar 30 08:29:15 2016, tara, DailyProgress, PMC, test on new PMC IMG_2528.JPGIMG_2540.JPG
                Reply  Wed Mar 30 10:28:44 2016, Evan, DailyProgress, PMC, test on new PMC 
                   Reply  Thu Mar 31 15:11:27 2016, tara, DailyProgress, PMC, test on new PMC 
                      Reply  Tue Apr 5 16:32:36 2016, tara, DailyProgress, PMC, test on new PMC 
                         Reply  Wed Apr 6 15:02:50 2016, tara, DailyProgress, PMC, test on new PMC 
                            Reply  Fri Apr 8 09:44:49 2016, tara, DailyProgress, PMC, test on new PMC IMG_2634.JPGIMG_2635.JPGIMG_2633.JPG
Message ID: 1617     Entry time: Wed Mar 30 08:29:15 2016     In reply to: 1616     Reply to this: 1619
Author: tara 
Type: DailyProgress 
Category: PMC 
Subject: test on new PMC 

We tested one of the two new PMCs. finesse is ~ 500 and the PZT works fine.

We assembled two new PMCs for testing how the new designed PMC would be. They looked great. The PZT wires were connected to BNC connectors.

We temporarily put the new PMC in the north path. The Faraday isolator (as an extra back reflection protection) was removed, and replaced by the new PMC. (The original marking for the PMC is still visible on the table, where the waist radius is 370 um). We used a post to clamped the PMC on the 2" block.

As a quick test, we have not mode matched the beam into the PMC yet. We just maximized the coupling by using two steering mirrors in front of the PMC.

Finesse measurement

  • Once we got TEM00 mode coming out, we scanned the laser and mesaured the transmission power.
  • The measured FWHM was 2.4 +/- 0.1 MHz. (The calibration for laser fast mod was taken from PSL:182)
  •  The FSR of the cavity = c/L ~ 1.2 GHz. So the Finesse = FSR/FWHM ~ 500 +/- 20.  

PZT test

  • We checked if the PZT works or not by applying high voltage (0-100 V) and see how many resonance of TEM00 we could find.
  • The slow offset (0.275 V) was used to set the first resonance at 0V, then the other resonance occrued at (16 , 32 , 41, 55, 71 V). They were not very linear, but it was likely to be the drift of the laser. And we were turning the knob bak and forth to find the resonance, so there might be some hysteriasis like behavior here. I will check it again.
  • According to the datasheet(add link), the pzt range is 3.3 um, with 200 V input. We should not have a problem with locking both PMC and refcav simultaneously anymore. By applying ~ 100 V, we can shift the PMC by 5 FSR (~ 6GHz), we can expect it to cover ~ 10 FSR (12 GHz) for the range of 200 V and refcav's FSR is 4GHz apart. So The PMC whole range (10 GHz) should be able cover at least 2 refcav modes (8GHz).

To Do:

  • Antonio will design a permanent mount for the new generic PMC, and check HOM in order to determine the sideband frequency we can use for locking.
  • We will try to lock the new PMC using the existing equipment, to determine loss and see if we can bring the PMC system back.
  • Test the 2nd PMC cavity.
  • We may fix the glass PMC, as Evan found out that it was broken long time ago, see PSL:1182
  • I need to look up the numbers for PMC (cavity length, mirror's reflectivity, ROC, designed finesse, spotsize, for comparing with the measurement and mode matching).
  • If time permits, assemble and characterize the rest of the PMCs (might be a project for a SURF student?). 
ELOG V3.1.3-